OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -14, and 5 + 8i f(x) = x4 - 362.5x2 + 1450x - 4984 f(x) = x4 - 9x3 + 32x2 - 725x + 4984 f(x) = x4 - 67x2 + 1450x - 4984 f(x) = x4 - 9x3 - 32x2 + 725x - 4984

OpenStudy (anonymous):

you know it will look like \((x-4)(x+14)(some quadratic)\)

OpenStudy (anonymous):

the hard part is finding the quadratic with zeros \(5+8i\) and \(5-8i\) do you know how to do that?

OpenStudy (anonymous):

i can show you two quick ways to do it if you like one is pretty quick, the other is amazingly quick, although you have to memorize something

OpenStudy (anonymous):

I'd love a demonstration, thanks!

OpenStudy (anonymous):

ok one way is to work backwards put \[x=4+8i\] subtract 4 get \[x-4=8i\] square (carefully) and get \[(x-4)^2=-64\] or \[x^2-8x+16=-64\] then add \(64\) and get \[x^2-8x+80=0\] and your quadratic is \(x^2-8x+80\)

OpenStudy (anonymous):

ahhh damn damn damn, it was \(5+8i\) not \(4+8i\)

OpenStudy (anonymous):

no matter, we can adjust quickly \[x=5+8i\\ x-5=8i\\ (x-5)^2=-64\\ x^2-10x+25=-64\\ x^2-10x+89=0\]

OpenStudy (anonymous):

really quick method requires memorizing that if \(a+bi\) is a zero of a quadratic, the quadratic is \[x^2-2ax+(a^2+b^2)\] in this case \(a=5, b=8\) so the quadratic is \[x^2-2\times 5x+(5^2+8^2)=x^2-10x+89\]

OpenStudy (anonymous):

your last job is to multiply \[(x-4)(x+14)(x^2-10x+89)\] without making an algebra error good luck with that, i would cheat

OpenStudy (anonymous):

Wow, that was incredibly helpful. Thank you so much!

OpenStudy (anonymous):

yw